A 900 Kg car is traveling east at 15 m/s when it collides with a 750 Kg car traveling north at 20 m/s. The cars stick together. What is the kinetic energy of the system after the collision?
Answer #1

let the x-component of final speed be Vx and y-component be Vy  

By the conservation of momentum in x-direction we have

900*15 + 750*0 = (900+750)Vx

Vx = 8.18 m/s

Conserving the momentum in y-direction we have

900*0 + 750*20 = (900+750)Vy

Vy = 9.1 m/s

V^2 = Vx^2 + Vy^2 = 8.18^2 + 9.1^2 = 149.55

Kinetic energy = 0.5(m1 + m2)V^2

= 0.5*(900+750)*149.55 = 123384.5 J